OpenStudy (jenny12):

Find the height of a trapezoid when the area is 187 feet squared, b1 is 15 ft. and b2 is 19 ft.

OpenStudy (anonymous):

h(15+19)/2=187

OpenStudy (anonymous):

\[h(15+19)/2=187\]

OpenStudy (anonymous):

h=187/17

OpenStudy (anonymous):

Area of a trapezoid is given by the formula Area = h (sum of parallel sides) -------------------- 2

OpenStudy (anonymous):

h=11

OpenStudy (anonymous):

Does that help?

OpenStudy (anonymous):

Yes, Pi solved it for u